Phantomdex:

There is a 25% chance that a book has a red cover, a 32% chance that a book is fiction, and an 8% chance that a book is fiction with a red cover. Determine the probability that a book has a red cover or is fiction. 65% 57% 49% 8%

toga:

To determine the probability that a book has a red cover or is fiction, we can use the formula: P(Red or Fiction) = P(Red) + P(Fiction) - P(Red and Fiction) Using the information given in the problem, we can plug in the probabilities: P(Red or Fiction) = 0.25 + 0.32 - 0.08 P(Red or Fiction) = 0.49 Therefore, the probability that a book has a red cover or is fiction is 49%, which is option C.